# Data-Efficient Learning for Generalizable Surgical Video Understanding

### v1

@@
11institutetext: Department of Information Technology (ITEC), University of Klagenfurt
11email: Sahar.Nasirihaghighi@aau.at
Data-Efficient Learning for Generalizable Surgical Video Understanding
Sahar Nasirihaghighi
1 Abstract
@@
2 Main Research Report
2.1 Research Problem and Motivation
@@ 2.2 Relevance and Related Work
Understanding surgical workflows through video analysis is a foundational task in computer-assisted interventions, enabling both intraoperative guidance and postoperative assessment. Initial approaches typically used image-level deep models that analyzed frames independently, but these failed to capture the temporal continuity of surgical procedures. To address this limitation, hybrid architectures combining spatial encoders with temporal sequence models such as CNN-RNNs were introduced, significantly improving performance [8, 12]. For example, TeCNO by Czempiel et al. [2] uses a multi-stage Temporal Convolutional Network (TCN) to enhance temporal consistency and phase recognition.
Although supervised methods perform well in controlled settings, their generalization across surgical domains is limited due to reliance on large annotated datasets and sensitivity to domain shifts across institutions or procedures. To address these limitations, semi-supervised learning (SSL) has gained traction by leveraging large amounts of unlabeled surgical data alongside limited annotations. Among the explored strategies, three main paradigms have shown particular promise: contrastive learning [24, 18], consistency regularization [13, 27], and pseudo-supervision [4, 28]. These approaches help reduce annotation costs while maintaining strong performance in surgical phase recognition.
@@ 2.3 Scientific Approach
Our early work focused on fine-grained action recognition in gynecologic laparoscopic surgery, a domain marked by complex visual conditions such as smoke, bleeding, occlusion, and rapid camera motion. We proposed a CNN-RNN hybrid architecture that integrates spatial feature extractors with stacked recurrent layers to capture temporal dependencies critical for recognizing short and variable actions [22]. To improve robustness to surgeon-specific variability such as differences in speed, technique, and temporal execution, we introduced a frame sampling strategy that enhances temporal diversity during training as illustrated in figure 2. Building on this foundation, we addressed the detection of critical intraoperative events such as bleeding. Although gynecologic laparoscopy is minimally invasive, intraoperative bleeding remains a safety-critical event that can affect surgical outcomes. To better model long-range temporal dependencies, we extended our earlier architecture by integrating transformer encoders into a hybrid CNN-transformer framework [20]. Spatial features extracted by were passed through multi-head self-attention layers, enabling the model to recognize temporally complex events including abdominal access, bleeding, needle passing, and coagulation or transection.
@@ 2.4 Proposed Solution
Only the top $50\%$ of consistent pseudo-labels are retained. To further filter unreliable predictions, DIST enforces two invariance constraints: temporal invariance (across different frame sampling strategies) and transformation invariance (under strong augmentations). Only labels satisfying both conditions are used to train a student model. In Stage 2, the student becomes the teacher, and the process is repeated. Evaluated on Cataract-1K and Cholec80, DIST outperforms both supervised and state-of-the-art SSL methods, especially in low-label scenarios. Its dual filtering strategy significantly improves pseudo-label robustness and generalization. In Stage 1, a teacher model generates pseudo-labels, which are filtered by a reliability estimation technique based on prediction consistency across three checkpoints
@@ 2.5 Results and Contributions
2.6 Open Challenges and Future Work
While my research has advanced supervised and semi-supervised approaches for surgical video analysis, important challenges remain for clinical deployment. A primary concern is domain shift, caused by variability in imaging systems, surgical protocols, and environments across institutions, which limits model generalization and reliability in unseen settings. As part of my final research objective, I am investigating strategies such as domain adaptation and semi-supervised federated learning to enhance robustness without relying on centralized data sharing. Semi-supervised federated learning is particularly valuable in real-world healthcare scenarios, where labeled data is scarce and strict privacy regulations prohibit the transfer of patient data. This approach enables decentralized training across multiple hospitals by leveraging abundant local unlabeled data alongside a small shared labeled dataset. It not only preserves patient confidentiality but also improves generalization across heterogeneous data sources. Moreover, it supports continuous learning on evolving clinical data, making it a practical solution for scalable and privacy-aware deployment in surgical AI systems.
2.7 Dissertation Status and Long-Term Goals
I am currently in the final stage of my PhD and plan to defend my dissertation in the near future. Over the past years, I have developed a comprehensive research pipeline aimed at advancing surgical video understanding through supervised and semi-supervised learning across diverse domains, including gynecologic laparoscopy and cataract surgery. My long-term vision is to contribute to the development of clinically viable, generalizable, and real-time AI systems that support surgical decision-making, improve patient safety, and enhance surgical training. I intend to pursue an academic or research-oriented career, continuing to bridge the gap between deep learning and clinical applications, fostering interdisciplinary collaboration, and leading impactful research at the intersection of AI and surgery.

### v2

@@
Abstract
@@
1 Research Problem and Motivation
@@ 2 Background
Understanding surgical workflows through video analysis is a foundational task in computer-assisted interventions, enabling both intraoperative guidance and postoperative assessment. Initial approaches typically used image-level deep models that analyzed frames independently, but these failed to capture the temporal continuity of surgical procedures. To address this limitation, hybrid architectures combining spatial encoders with temporal sequence models such as CNN-RNNs were introduced, significantly improving performance [8]. For example, TeCNO by Czempiel et al. [2] uses a multi-stage Temporal Convolutional Network (TCN) to enhance temporal consistency and phase recognition. Although supervised methods perform well in controlled settings, their generalization across surgical domains is limited due to reliance on large annotated datasets and sensitivity to domain shifts across institutions or procedures. To address these limitations, semi-supervised learning (SSL) has gained traction by leveraging large amounts of unlabeled surgical data alongside limited annotations. Among the explored strategies, three main paradigms have shown particular promise: contrastive learning [23, 17], consistency regularization [12, 26], and pseudo-supervision [4, 27]. These approaches help reduce annotation costs while maintaining strong performance in surgical phase recognition.
@@ 3 Scientific Approach
Our early work focused on fine-grained action recognition in gynecologic laparoscopic surgery, a domain marked by complex visual conditions such as smoke, bleeding, occlusion, and rapid camera motion. We proposed a CNN-RNN hybrid architecture that integrates spatial feature extractors with stacked recurrent layers to capture temporal dependencies critical for recognizing short and variable actions [21]. To improve robustness to surgeon-specific variability such as differences in speed, technique, and temporal execution, we introduced a frame sampling strategy that enhances temporal diversity during training as illustrated in Fig. 2. Building on this foundation, we addressed the detection of critical intraoperative events such as bleeding. Although gynecologic laparoscopy is minimally invasive, intraoperative bleeding remains a safety-critical event that can affect surgical outcomes. To better model long-range temporal dependencies, we extended our earlier architecture by integrating transformer encoders into a hybrid CNN-transformer framework [19]. Spatial features extracted by were passed through multi-head self-attention layers, enabling the model to recognize temporally complex events including abdominal access, bleeding, needle passing, and coagulation or transection.
@@ 4 Proposed Solution
Only the top $50\%$ of consistent pseudo-labels are retained. To further filter unreliable predictions, DIST enforces two invariance constraints: temporal invariance (across different frame sampling strategies) and transformation invariance (under strong augmentations). Only labels satisfying both conditions are used to train a student model. In Stage 2, the student becomes the teacher, and the process is repeated. Evaluated on Cataract-1K and Cholec80, DIST outperforms both supervised and state-of-the-art SSL methods, especially in low-label scenarios. Its dual filtering strategy significantly improves pseudo-label robustness and generalization. In Stage 1, a teacher model generates pseudo-labels, which are filtered by a reliability estimation technique based on prediction consistency across three checkpoints.
@@ 5 Results and Contribution
6 Open Challenges and Future Work
While my research has advanced supervised and semi-supervised surgical video analysis, clinical deployment challenges remain. Domain shift–caused by variability in imaging systems, protocols, and environments across institutions–limits model generalization in unseen settings. I am investigating domain adaptation and federated learning to enhance robustness without centralized data sharing. Federated semi-supervised learning is particularly valuable in healthcare, where labeled data is scarce and privacy regulations prohibit patient data transfer. This approach enables decentralized training across hospitals by leveraging local unlabeled data alongside small shared labeled datasets. It preserves patient confidentiality while improving generalization across heterogeneous sources and supports continuous learning on evolving clinical data, providing a practical solution for scalable, privacy-aware deployment in surgical AI systems.
7 Long Term Goals
My research focuses on advancing surgical video understanding through supervised and semi-supervised learning across gynecologic laparoscopy and cataract surgery. My vision is to develop clinically deployable, generalizable, real-time AI systems that strengthen surgical decision-making, improve patient safety, and enhance surgical training. I intend to pursue an academic research career bridging deep learning and clinical practice through interdisciplinary collaboration, translating technological innovations into tangible clinical impact.